# Data $\times$ LLM: From Principles to Practices

### v2

@@ Abstract
The integration of large language model (LLM) and data management (DATA) is rapidly redefining both domains. In this survey, we comprehensively review the bidirectional relationships. On the one hand, DATA4LLM, spanning large-scale data processing, storage, and serving, feeds LLMs with high quality, diversity, and timeliness of data required for stages like pre-training, post-training, retrieval-augmented generation, and agentic workflows: $(i)$ Data processing for LLMs includes scalable acquisition, deduplication, filtering, selection, domain mixing, and synthetic augmentation; $(ii)$ Data Storage for LLMs focuses on efficient data and model formats, distributed and heterogeneous storage hierarchies, KV-cache management, and fault-tolerant checkpointing; $(iii)$ Data serving for LLMs tackles challenges in RAG (e.g., knowledge post-processing), LLM inference (e.g., prompt compression, data provenance), and training strategies (e.g., data packing and shuffling). On the other hand, in LLM4DATA, LLMs are emerging as general-purpose engines for data management. We review recent advances in $(i)$ data manipulation, including automatic data cleaning, integration, discovery; $(ii)$ data analysis, covering reasoning over structured, semi-structured, and unstructured data, and $(iii)$ system optimization (e.g., configuration tuning, query rewriting, anomaly diagnosis), powered by LLM techniques like retrieval-augmented prompting, task-specialized fine-tuning, and multi-agent collaboration.
@@ 1 INTRODUCTION
Example-① Data Processing for LLMs. Processing a large-scale training dataset (e.g., $\sim$4 TB multi-modal tokens utilized in Qwen2.5-VL pretraining [70]) poses several challenges. First, acquiring diverse raw data (e.g., over 10,000 object categories for visual grounding) demands substantial efforts in data collection (Section 2.2.1) and, in many cases, data synthesis (Section 2.2.6). Second, preparing high-quality training samples requires robust preprocessing, including rigorous data filtering (Section 2.2.3), along with dedicated evaluation approaches. Third, the overall performance of LLMs depends heavily on an end-to-end pipeline that effectively schedules and coordinates these processing tasks, especially for the pretraining stage (Section 2.2.7).
@@
Example–③ Data Serving for LLMs. Data serving plays a critical role in selecting and preparing input data (e.g., the task-specific prompts), directly affecting the quality of LLM’s responses. Taking retrieval-augmented generation (RAG) as an example, EyeLevel.ai [37] observed that when relying solely on vector similarity, RAG accuracy declines notably with 10,000-page documents, and the performance degradation can reach up to 12% with 100,000 pages (still fewer than enterprise-scale datasets). Several challenges arise in this context. First, the retrieved knowledge is typically noisy and must be filtered and re-ranked to ensure relevance and factual accuracy (Section 2.4.1). Second, the retrieved content is often lengthy exceeding the input capacity or comprehension of LLMs, necessitating effective compression techniques to preserve utility while improving performance (Section 2.4.2).
@@ 1.1 Techniques of DATA4LLM
Data Characteristics across LLM Stages (§ 2.1). As shown in Figure 1, datasets play a critical role in enabling the desired capabilities at each LLM stage, including (1) pre-training, (2) continual pre-training, (3) fine-tuning, (4) reinforcement learning, (5) retrieval-augmented generation (RAG), (6) LLM agents, and (7) evaluation. For each stage, we separately analyze the required data across multiple dimensions (e.g., data format, scale, and domain diversity).
@@ 1.3 Comparison with Existing Surveys
Different from existing LLM and data management surveys [413, 55, 86, 405, 274, 276, 380, 496], our survey offers a comprehensive and detailed overview of the key intersections between LLMs and data management, highlighting how they can mutually benefit from each other. We uniquely position our work at the intersection of data for LLMs (e.g., how to acquire, process, store, and serve data for training and using LLMs) and LLMs for data (e.g., how LLMs can be leveraged to enhance data management tasks).
$\bullet$ We investigate the unique characteristics of data across different LLM development stages (Figure 2), and provide a systematic overview of the associated challenges and techniques in data processing, storage, and serving (Table I). Instead, prior surveys [413, 55, 86] primarily center on the pre-training stage without covering the full LLM lifecycle like supervised fine-tuning (SFT), retrieval-augmented generation (RAG), and agent-based applications.
@@ 2 Data Management for LLM (DATA4LLM)
2.1 Data Characteristics across LLM Stages
@@
Data Storage KV Cache N/A N/A N/A Cache Space Managemen KV Indexing KV Placement KV Shrinking KV Placement KV Shrinking N/A
@@
Figure 3: Example LLM Data Distribution - (a) General Domain (SFT)[111], (b) General Domain (Eval) [246], (c) Law (SFT)[456], (d) Law (Eval)[116], (e) Code (SFT) [296], (f) Code (Eval)[210].
Before discussing specific data management techniques for LLMs, we first provide an overview of the data characteristics across different stages of LLM development. These characteristics include (1) the data scale, (2) the drawn data domain diversity, and (3) the data formats (e.g., question-answer pairs with or without explicit labels). Given the differences in these characteristics across stages, distinct techniques for data processing, storage, and serving are required (Table I).
Data for Pretraining. In the pre-training stage, LLMs rely on TB-scale, diverse datasets to acquire broad language and even cross-modality understanding capabilities, while reducing the risk of overfitting. These datasets are typically sourced from a wide range of domains and formats, including web crawls (e.g., HTML pages and WARC files) [11], open-source code repositories (e.g., raw source code files with metadata) [14], books (e.g., plain text or EPUB formats) [505], academic papers (e.g., LaTeX source or PDF-converted text)  [2], and interleaved image-text corpora (e.g., aligned captioned images in JSON or WebDataset format) [226].
@@ 2.2.5 Data Mixing
Heuristic Optimization provides empirical data mixing strategies such as setting different ratios of datasets based on various factors (e.g., complexity and diversity of the datasets) that likely improve LLMs’ abilities.
Bilevel Optimization. Bilevel Optimization [109] is a closed-loop optimization technique that ensures model parameters are well optimized. Generally, Bilevel optimization involves two nested optimization problems: (1) the inner-level problem ensures model parameters are optimized under given weights (e.g., minimizing weighted training loss), while (2) the outer-level updates weights through backpropagation of validation loss, forming a closed-loop optimization.
TABLE VI: Comparison of Data Mixing Methods for LLM.
@@
Heuristic Optimization Pre-training Multi-Source Data Adjusting Intuitive and easy to implement, suitable for rapid experimentation.
@@
Bilevel Optimization Pre-training Calculate domain contribution by Requires a proxy model, performances well in OOD datasets.
gradient inner products[136]
Distributionally Robust Optimization Pre-training Group DRO [429] Only need one proxy model compared to [280]
which uses a proxy model and a reference model
Fine-tuning Task-level DRO [280] Perform well with problems with significant
loss difference across datasets
Model-Based Optimization Pre-training Linear Regression Model [265] Only 10% of DoReMi’s [429] computational resources are required.
@@
Distributionally Robust Optimization. To search for a robust data mixing strategy (which can be sub-optimal but with low uncertainty), some methods adopt Distributionally Robust Optimization (DRO) for data mixing. DRO achieves robustness against distributional uncertainty by optimizing for the worst-case scenario within a set of distributions (referred to as the uncertainty set or ambiguity set).
Model-Based Optimization. This category of methods design linear or non-linear models that depict $(i)$ the relation between the distribution of each domain, $(ii)$ validation loss, and $(iii)$ some other variables like training steps, based on which they find the optimal settings through various model-based techniques.
@@
Typically, REGMIX [265] defines the domains by source (like ArXiv, FreeLaw, etc.), which uses Dirichlet distribution (which controls the distribution of probabilities across multiple categories with a parameter) to generate all kinds of data distribution of several domains to trains small-scale proxy model to collect performance data, which is then used to fit a linear regression model (LightGBM [207]) to predict the optimal data mixing distribution. Then REGMIX uses both the best distribution and the average of top-100 distributions to verify on variations of TinyLlama [468] with additional layers with versions of 1B and 7B.
@@
$\bullet$ Exponential Functions. [448] establishes an exponential relationship between validation loss and data mixing ratios of several domains (e.g., public datasets like Pile-CC, Books3), $L({r})=c+k\exp\left(\sum_{i}t_{i}r_{i}\right)$, where $L({r})$ is the validation loss, ${r}$ represents the mixing ratios of different domains, and $c$, $k$, and $t_{i}$ are learnable parameters. That is, it experiments on a small model with the exponential relationships to predict the best data domain mixing ratios on LLM performance with scaling laws, which combines training step scaling laws ($L(S)=c+kS^{\alpha}$, where $S$ is the number of training steps, and $\alpha$ is a fitting parameter.), which is used to infer the validation loss at target training steps from results at smaller steps, and model size scaling laws ($L(N)=c+kN^{\beta}$, where $N$ is the number of model parameters, and $\beta$ is a fitting parameter), which is used to infer the validation loss for large model sizes from smaller model sizes.
$\bullet$ Classification Model. [253] aims to find the data proportion of closed-source model by data proportion detection, which first generating large-scale data from the LLM, then using a classification model to categorize the generated data and compute perplexity, deriving the proportions of pre-training data based on the Data Mixing Law (which is a mathematical formula describing the relationship between the proportion of pre-training data and the model’s loss in different domains.).
@@ 2.2.6 Data Distillation and Synthesis
$\bullet$ Task-Specific Prompt Distillation. To significantly reduce inference costs and latency while maintaining performance, [shirgaonkar202x4knowledge] employs task-specific prompts: (1) Chain-of-Density (CoD): Iteratively adds entities to summarize for enhanced density. (2) Chain-of-Thought (CoT): Guides reasoning tasks (e.g., math) through stepwise logic. Using GSM8K [107] data and Llama-3.1-405B-Instruct, synthetic data is generated for fine-tuning smaller models (Llama-3.1-8B/70B-Instruct) paired with simplified prompts, balancing efficiency and task specialization.
@@ 2.3.3 Data Organization
LeanVec [386] combines linear dimensionality reduction with LVQ for vector compression. In ID (In-distribution) scenarios, LeanVec uses PCA, while in OOD (Out-of-distribution) scenarios, it introduces the LeanVec-OOD optimization method, which identifies the optimal projection subspace for both the dataset and the query set by minimizing the squared inner product between the query vector and the representation error, thereby reducing vector dimensionality. However, LeanVec is a simple linear dimensionality reduction method, and its performance may be affected in terms of accuracy when reducing the dimensionality drastically.
@@ 2.3.4 Data Movement
SiloD [475] leverages the characteristics of the pipelined execution of data loading and computation at the pre-training and SFT stage to build an enhanced performance evaluator. When data loading becomes the bottleneck, it uses a learned model (IOPerf) to quantify the cache and remote I/O demands of different training jobs,providing support for resource allocation in the pipelined execution of data loading and computation.
@@ 2.4.1 Data Shuffling
Data Selection for RAG. In the RAG stage, it is necessary to retrieve the stored knowledge (see details in 2.3.3) and provided the retrieved results to the LLM. During this process, it needs to ensure the effectiveness of the retrieved results in order to obtain better answers from the LLM [282]. Currently, the retrieval quality is mainly guaranteed through RAG knowledge filtering and RAG knowledge re-ranking.

### v3

@@ Abstract
The integration of large language model (LLM) and data management (DATA) is rapidly redefining both domains. In this survey, we comprehensively review the bidirectional relationships. On the one hand, DATA4LLM, spanning large-scale data processing, storage, and serving, feeds LLMs with diversity, redundant, high quality, and sanitized data (following the “IaaS” concept) required for stages like pre-training, post-training, retrieval-augmented generation, and agentic workflows: $(i)$ Data processing for LLMs includes scalable acquisition, deduplication, filtering, selection, domain mixing, and synthetic augmentation; $(ii)$ Data Storage for LLMs focuses on efficient data and model formats, distributed and heterogeneous storage hierarchies, KV-cache management, and fault-tolerant checkpointing; $(iii)$ Data serving for LLMs tackles challenges in RAG (e.g., knowledge post-processing), LLM inference (e.g., prompt compression, data provenance), and training strategies (e.g., data packing and shuffling). On the other hand, in LLM4DATA, LLMs are emerging as general-purpose engines for data management. We review recent advances in $(i)$ data manipulation, including automatic data cleaning, integration, discovery; $(ii)$ data analysis, covering reasoning over structured, semi-structured, and unstructured data, and $(iii)$ system optimization (e.g., configuration tuning, query rewriting, anomaly diagnosis), powered by LLM techniques like retrieval-augmented prompting, task-specialized fine-tuning, and multi-agent collaboration.
@@ 1 INTRODUCTION
Example-① Data Processing for LLMs. Processing a large-scale training dataset (e.g., $\sim$4 TB multi-modal tokens utilized in Qwen2.5-VL pretraining [70]) poses several challenges. First, acquiring diverse raw data (e.g., over 10,000 object categories for visual grounding) demands substantial efforts in data collection (Section 2.3.1) and, in many cases, data synthesis (Section 2.3.6). Second, preparing high-quality training samples requires robust pre-processing, including rigorous data filtering (Section 2.3.3), along with dedicated evaluation approaches. Third, the overall performance of LLMs depends heavily on an end-to-end pipeline that effectively schedules and coordinates these processing tasks, especially for the pretraining stage (Section 2.3.7).
@@
Example–③ Data Serving for LLMs. Data serving plays a critical role in selecting and preparing input data (e.g., the task-specific prompts), directly affecting the quality of LLM’s responses. Taking retrieval-augmented generation (RAG) as an example, EyeLevel.ai [37] observed that when relying solely on vector similarity, RAG accuracy declines notably with 10,000-page documents, and the performance degradation can reach up to 12% with 100,000 pages (still fewer than enterprise-scale datasets). Several challenges arise in this context. First, the retrieved knowledge is typically noisy and must be filtered and re-ranked to ensure relevance and factual accuracy (Section 2.5.1). Second, the retrieved content is often lengthy and exceeds the input capacity or comprehension of LLMs, necessitating effective compression techniques to preserve utility while improving performance (Section 2.5.2).
@@ 1.1 Techniques of DATA4LLM
Characteristics of LLM Datasets (§ 2.2). As shown in Figure 1, datasets (following the “IaaS” concept) play a critical role in enabling the desired capabilities at each LLM stage, including (1) pre-training, (2) continual pre-training, (3) fine-tuning, (4) reinforcement learning, (5) retrieval-augmented generation (RAG), (6) LLM agents, and (7) evaluation. For each stage, we separately analyze the characters of required data (e.g., preferred formats and emphasized aspects within IaaS) and the corresponding data techniques (see Table I).
@@ 1.3 Comparison with Existing Surveys
Different from existing LLM and data management surveys [405, 55, 86, 398, 272, 274, 374, 488], our survey offers a comprehensive and detailed overview of the key intersections between LLMs and data management, highlighting how they can mutually benefit from each other. We uniquely position our work at the intersection of data for LLMs (e.g., how to acquire, process, store, and serve LLM data) and LLMs for data (e.g., how LLMs can be leveraged to enhance data management tasks).
$\bullet$ We propose the IaaS concept as a principled lens to assess LLM dataset quality. The IaaS concept identifies four essential dimensions, including inclusiveness, abundance, articulation, and sanitization. This concept is promising to offers an evaluative criteria for guiding data management and understanding its impact across the LLM development lifecycle (see Section 2.1).
$\bullet$ We investigate the unique characteristics of data across different LLM development stages (Figure 2), and provide a systematic overview of the associated challenges and techniques in data processing, storage, and serving (Table I). In contrast, prior surveys [405, 55, 86] primarily center on the pre-training stage without covering the full LLM lifecycle like supervised fine-tuning (SFT), retrieval-augmented generation (RAG), and agent-based applications.
@@ 2 Data Management for LLM (DATA4LLM)
2.1 IaaS” Concept of LLM Data
Based on our investigation of over 400 papers  [^2] https://github.com/weAIDB/awesome-data-llm, we introduce the IaaS concept for evaluating the quality of LLM datasets.
@@ 2.1 IaaS” Concept of LLM Data
Data Storage KV Cache N/A N/A N/A Cache Space Management KV Indexing KV Placement KV Shrinking KV Placement KV Shrinking N/A
@@
Figure 3: Example LLM Data Distributions - (a) General Domain (SFT)[110], (b) General Domain (Eval) [244], (c) Law (SFT)[447], (d) Law (Eval)[115], (e) Code (SFT) [294], (f) Code (Eval)[208].
(1) Inclusiveness: LLMs require data with broad and diverse coverage across multiple dimensions, including domains (e.g., general knowledge, specialized fields like finance, medicine, math [98], and physics [233]), task types (e.g., question answering, summarization, code completion [401, 290, 353, 45, 436]), data sources (e.g., GitHub, Wikipedia [149, 11, 330, 347]), languages [93, 347], expression styles (e.g., academic, casual, formal [282, 470]), and data modalities (e.g., text [149, 11], images [145, 185], videos [437, 216, 390], tables [330]).
(2) abundance: LLMs require data with appropriate volume and balanced composition to prevent overfitting on homogeneous data. Specifically, abundance of data involves: $(i)$ constructing well-balanced datasets during pre-training [139, 302, 420, 263], $(ii)$ adjusting data ratios to align with target applications during fine-tuning [278, 135], and $(iii)$ continually enhancing domain-specific capabilities while maintaining acceptable general performance degradation in continual pre-training [323, 160]. Notably, the strength of LLMs lies not only in large-scale data [282, 481, 11, 330, 149, 347], but also in constructing purposefully balanced datasets, which can further accelerate training and reduce computational cost.
(3) articulation: LLMs require data that exhibit strong articulation, including three key aspects: $(i)$ the data should be well-formatted (e.g., proper punctuation and capitalization [90]), clean (free from duplicates, typos, and irrelevant content such as spam or gibberish [90]), and self-contained, featuring clear, fluent, and unambiguous language [282, 470], $(ii)$ the data should be instructive [178, 179, 98], i.e., offering sufficient context, guidance, and intermediate explanations that help the model connect questions to relevant background knowledge and understand the reasoning process. $(iii)$ the data should involve step-by-step reasoning[230, 442, 346, 173, 496], such that enhancing the LLMs’ reasoning capabilities by decomposing complex tasks into smaller, interpretable steps.
(4) Sanitization: LLMs require data to be sanitized, meaning it is rigorously controlled and filtered to remove harmful elements while maintaining inclusiveness and neutrality. This involves four critical dimensions: $(i)$ Privacy compliance, which requires the exclusion of personally identifiable information (e.g., ID numbers, phone numbers), inferred social relationships, and geolocation-related metadata [450, 268, 275]; $(ii)$ Toxicity-free content, ensuring the complete removal of hate speech, incitement to violence, and psychologically harmful language, as well as eliminating any discriminatory or aggressive semantic constructs [296]; $(iii)$ Ethical consistency, which prohibits the presence of extremist ideologies, instructions for illegal activities, and stereotype-reinforcing narratives that may cause social harm [345, 360, 296]; and $(iv)$ Risk mitigation, filtering out unverified medical claims, politically misleading information, and culturally insensitive expressions to prevent misinformation and value misalignment. Sanitized data must maintain a neutral tone and adopt an inclusive contextual framework, serving as a critical foundation for building safe LLMs [345, 360].
2.2 Data Characters across LLM Stages
Next we specifically discuss the data characteristics across different LLM stages, together with the distinct techniques for data processing, storage, and serving (Table I).
Data for Pretraining. In the pre-training stage, LLMs rely on TB-scale, diverse datasets to acquire broad language and even cross-modality understanding capabilities, while reducing the risk of overfitting. These datasets are typically sourced from a wide range of domains and formats, including web crawls (e.g., HTML pages and WARC files [11]), open-source code repositories (e.g., raw source code files with metadata [14]), books (e.g., plain text or EPUB formats [497]), academic papers (e.g., LaTeX source or PDF-converted text [2]), and interleaved image-text corpora (e.g., aligned captioned images in JSON or WebDataset format [224]).
@@ 2.3.5 Data Mixing
TABLE VI: Comparison of Data Mixing Methods for LLMs.
@@
Before Training (Human Experience) Pre-training Multi-Source Data Adjusting Intuitive and easy to implement, suitable for rapid experimentation.
@@
Before Training (Model-Based Optimization) Pre-training Linear Regression Model [263] Only 10% of DoReMi’s [420] computational resources are required.
@@
During Training (Bilevel Optimization) Pre-training Calculate domain contribution by Requires a proxy model, performances well in OOD datasets.
gradient inner products[135]
During Training (Distributionally Robust Optimization) Pre-training Group DRO [420] For pre-training, smooth adjusting to prevent abrupt weight changes
Fine-tuning Task-level DRO [278] For fine tuning, quick response to task difficulty changes
Before-Training Mixing (Human Experience). This method provides empirical data mixing strategies such as setting different ratios of datasets based on various factors (e.g., complexity and diversity of the datasets) that likely improve LLMs’ abilities.
Before-Training Mixing (Model-Based Optimization). This category of methods design linear or non-linear models that depict $(i)$ the relation between the distribution of each domain, $(ii)$ validation loss, and $(iii)$ some other variables like training steps, based on which they find the optimal settings through various model-based techniques.
@@
Typically, REGMIX [263] defines the domains by source (like ArXiv, FreeLaw, etc.), which uses Dirichlet distribution (which controls the distribution of probabilities across multiple categories with a parameter) to generate all kinds of data distribution of several domains to train a small-scale proxy model to collect performance data, which is then used to fit a linear regression model (LightGBM [205]) to predict the optimal data mixing distribution. Then REGMIX uses both the best distribution and the average of top-100 distributions to verify on variations of TinyLlama [459] with additional layers with versions of 1B and 7B.
@@
$\bullet$ Exponential Functions. Data Mixing Law [439] establishes an exponential relationship between validation loss and data mixing ratios of several domains (e.g., public datasets like Pile-CC, Books3), $L({r})=c+k\exp\left(\sum_{i}t_{i}r_{i}\right)$, where $L({r})$ is the validation loss, ${r}$ represents the mixing ratios of different domains, and $c$, $k$, and $t_{i}$ are learnable parameters. That is, it experiments on a small model with the exponential relationships to predict the best data domain mixing ratios on LLM performance with scaling laws, which combines training step scaling laws ($L(S)=c+kS^{\alpha}$, where $S$ is the number of training steps, and $\alpha$ is a fitting parameter.), which is used to infer the validation loss at target training steps from results at smaller steps, and model size scaling laws ($L(N)=c+kN^{\beta}$, where $N$ is the number of model parameters, and $\beta$ is a fitting parameter), which is used to infer the validation loss for large model sizes from smaller model sizes.
$\bullet$ Classification Model. [251] aims to find the data proportion of closed-source model by data proportion detection, which first generating large-scale data from the LLM, then using a classification model to categorize the generated data and compute perplexity, deriving the proportions of pre-training data based on the Data Mixing Law [439] (which is a mathematical formula describing the relationship between the proportion of pre-training data and the model’s loss in different domains.).
@@
During-Training Mixing (Bilevel Optimization). This method adopts a closed-loop optimization technique that ensures model parameters are well optimized [108]. Generally, Bilevel optimization involves two nested optimization problems: (1) the inner-level problem ensures model parameters are optimized under given weights (e.g., minimizing weighted training loss), while (2) the outer-level updates weights through backpropagation of validation loss, forming a closed-loop optimization.
During-Training Mixing (Distributionally Robust Optimization). To search for a robust data mixing strategy (which can be sub-optimal but with low uncertainty), some methods adopt Distributionally Robust Optimization (DRO) for data mixing. DRO achieves robustness against distributional uncertainty by optimizing for the worst-case scenario within a set of distributions (referred to as the uncertainty set or ambiguity set).
@@ 2.3.6 Data Distillation and Synthesis
$\bullet$ Task-Specific Prompt Distillation. To significantly reduce inference costs and latency while maintaining performance, [353] employs task-specific prompts: (1) Chain-of-Density (CoD): Iteratively adds entities to summarize for enhanced density. (2) Chain-of-Thought (CoT): Guides reasoning tasks (e.g., math) through stepwise logic. Using GSM8K [106] data and Llama-3.1-405B-Instruct, synthetic data is generated for fine-tuning smaller models (Llama-3.1-8B/70B-Instruct) paired with simplified prompts, balancing efficiency and task specialization.
@@ 2.4.3 Data Organization
LeanVec [380] combines linear dimensionality reduction with LVQ for vector compression. In ID(In-distribution) scenarios, LeanVec uses PCA, while in OOD(Out-of-distribution) scenarios, it introduces the LeanVec-OOD optimization method, which minimizes the square of the inner product between the query vector and the representation error to find the optimal projection subspace for both the dataset and the query set, thereby reducing the vector dimension. However, LeanVec is a simple linear dimensionality reduction method, and its performance may be affected in terms of accuracy when reducing the dimensionality drastically.
@@ 2.4.4 Data Movement
SiloD [466] leverages the characteristics of the pipelined execution of data loading and computation at the pre-training and SFT stage to build an enhanced performance evaluator. When data loading becomes the bottleneck, it uses a learned model (IOPerf) to quantify the cache and remote I/O demands of different training jobs,providing support for resource allocation in the pipelined of data loading and computation.
@@ 2.5.1 Data Shuffling
Data Selection for RAG. In the RAG stage, it is necessary to retrieve the stored knowledge (see details in 2.4.3) and provided the retrieved results to the LLM. During this process, it needs to ensure the effectiveness of the retrieved results in order to obtain better answers from the LLM [280]. Currently, the retrieval quality is mainly guaranteed through RAG knowledge filtering and RAG knowledge re-ranking.